Find the orthogonal canonical reduction of the quadratic form -x^2+y^2+z^2-4xy-4xz.
Also, find it's principal axes.
1
Expert's answer
2019-03-21T10:01:24-0400
"f(x, y, z)=-x^2+y^2+z^2-4xy-4xz"

The matrix of the quadratic form is


"A=\begin{pmatrix}
 -1 & -2 & -2 \\
 -2 & 1 & 0 \\ -2 & 0 &1
\end{pmatrix}"

The characteristic equation is


"\begin{vmatrix}
 -1-\lambda & -2 & -2 \\
 -2 & 1-\lambda & 0 \\ -2 & 0 & 1-\lambda
\end{vmatrix}=0"

Hence


"(1-\lambda)({\lambda}^2-9)=0""\lambda_1=1, \lambda_2=3, \lambda_3=-3"

Thus, the orthogonal canonical reduction is


"Q=\begin{pmatrix}
 x' & y' & z' \\
\end{pmatrix}*\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 3 & 0 \\ 0 & 0 &-3
\end{pmatrix}*\begin{pmatrix}
 x' \\
 y' \\ z'
\end{pmatrix}=(x')^2+3(y')^2-3(z')^2"

Find eigenvectors:


"\lambda_1=1""\begin{pmatrix}
 -2 & -2 & -2 \\
 -2 & 0 & 0 \\ -2 & 0 &0
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 1 \\ 0 & 0 &0
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""v=\begin{pmatrix}
 0 \\
 -1 \\ 1
\end{pmatrix}"

The principal axis is


"{1 \over \sqrt{2}}\begin{pmatrix}
 0 \\
 -1 \\ 1
\end{pmatrix}""\lambda_2=3""\begin{pmatrix}
 -4 & -2 & -2 \\
 -2 & -2 & 0 \\ -2 & 0 &-2
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""\begin{pmatrix}
 1 & 0 & 1 \\
 0 & 1 & -1 \\ 0 & 0 & 0
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""v=\begin{pmatrix}
 -1 \\
 1 \\ 1
\end{pmatrix}"

The principal axis is


"{1 \over \sqrt{3}}\begin{pmatrix}
 -1 \\
 1 \\ 1
\end{pmatrix}""\lambda_3=-3""\begin{pmatrix}
 2 & -2 & -2 \\
 -2 & 4 & 0 \\ -2 & 0 &4
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""\begin{pmatrix}
 1 & 0 & -2 \\
 0 & 1 & -1 \\ 0 & 0 & 0
\end{pmatrix}*\begin{pmatrix}
 v_1 \\
 v_2 \\ v_3
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\ 0
\end{pmatrix}""v=\begin{pmatrix}
 2 \\
 1 \\ 1
\end{pmatrix}"

The principal axis is


"{1 \over \sqrt{6}}\begin{pmatrix}
 2 \\
 1 \\ 1
\end{pmatrix}"